OpenStudy (anonymous):

The velocity of sound in air is given by the equation where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity of sound in air is 369 meters per second. Round to the nearest degree. A. 507ºC B. 6,535ºC C. 7,081ºC D. 67ºC

OpenStudy (anonymous):

269^2/400 -273 = t 26961/400 67.4025 D. 67 degrees C ???????????????????????????????????????

OpenStudy (ajprincess):

what is the equation?

OpenStudy (phi):

the original equation was probably \[ v= 20 \sqrt{273+t} \] solve for t. The first step is square both sides \[ v^2= 400(273+t) \] divide both sides by 400 \[ \frac{v^2}{400} = 273+t \] subtract 273 from both sides \[ t= \frac{v^2}{400}-273 \] now find the numerical value of t when v= 369 \[ t= \frac{369^2}{400}-273= 67.4 \]